OpenStudy (anonymous):

1. A tunnel is 50m wide and it has a maximum height of 40m. How high is the tunnel when it is 5m its base?

OpenStudy (anonymous):

(50/40)=(5/h) Muiltiply cross 50h=40*5 h=200/50 h=4